OpenStudy (anonymous):

Find the common difference in an arithmetic sequence in which a10-a20=70.

OpenStudy (anonymous):

is tht u on the pic?:o

OpenStudy (anonymous):

yes? can u please help me now

OpenStudy (anonymous):

if u were real

OpenStudy (anonymous):

what is u problem

OpenStudy (anonymous):

check out the link

OpenStudy (anonymous):

if u don't wanna help screw off

OpenStudy (anonymous):

everyones here to help

OpenStudy (anonymous):

lol. both of you seem rather like mindless trolls.

OpenStudy (anonymous):

i hate fake peoples lmao

OpenStudy (anonymous):

math_Hater_ u a$$hole

OpenStudy (anonymous):

lol how?

OpenStudy (anonymous):

can anybody actually help me

OpenStudy (anonymous):

so you're telling that diff. of 10th, 20th term is 70. So let 1st term = a, n1 = 10, n2 = 20, and 10th -20th = 70. If d is common difference - The 10th term = a + (10-1).d => a+9d And 20th term = a + (20-1)d => a + 19d Now 10th term - 20th term = 70. a+9d - (a+19d) = 70. 9d - 19d = 70 70 = -10d. d = -7. Common difference = -7 = ANS.

OpenStudy (anonymous):

abc is retarded to pretend to be cute and lie about his pic, rather shallow and retarded, and to use the word "a$$hole". and math_hater- is not-so-smart for bringing it up and picking a fight about it.

OpenStudy (anonymous):

and lucky is a genius.

OpenStudy (anonymous):

@abc123! your worked out sum is there. cheers :)

OpenStudy (anonymous):

thank you my man

OpenStudy (anonymous):

lol sorry freakin

OpenStudy (anonymous):

who gave this fake guy a medal?

OpenStudy (anonymous):

@Math_Hater- am nt faker.

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

not u lucky

OpenStudy (anonymous):

im talking about abc123